OpenStudy (anonymous):

r=2/3t + v, for t show work please?…

OpenStudy (saifoo.khan):

\[r = \frac23t + v\]

OpenStudy (saifoo.khan):

\[-v + r = \frac 23t\]

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (anonymous):

(3r-3v)/2 = t

OpenStudy (saifoo.khan):

\[\frac{3(-v+r)}{3}=t\]

OpenStudy (anonymous):

\[(3r-3v)\div2 = t\]

OpenStudy (anonymous):

nonononon, over 2, not three.